# Probabilistic Design of Multi-Dimensional Spatially-Coupled Codes

### v1

@@ I Introduction
The finite-length performance advantage of SC codes results from the additional degrees of freedom they offer the code designer via partitioning. Those additional degrees of freedom grow with the memory of the SC code. There are various results on high performance SC code designs with low memories in the literature [7, 11, 12]. Some of these designs are optimal with respect to the number of detrimental objects [13]; others demonstrated notable performance gains in data storage and transmission systems [14, 15]. Optimal solutions are notorious to achieve for SC codes with high memories because of the remarkable complexity growth. Therefore, a recent idea introduced probabilistic design of SC codes with high memories, which is based on the gradient-descent (GD) algorithm, to find locally-optimal solutions with respect to the number of detrimental objects [8].
@@ II Preliminaries
$\mathbf{H}_{\textup{SC}}$ is obtained from some binary matrix $\mathbf{H}^{\textup{g}}_{\textup{SC}}$ by replacing each nonzero (zero) entry in it by a circulant (zero) $z\times z$ matrix, $z\in\mathbb{N}$. Throughout the paper, the notation “g" in the superscript of any matrix refers to its protograph matrix, where each circulant matrix is replaced by $1$ and each zero matrix is replaced by 00. $\mathbf{H}^{\textup{g}}_{\textup{SC}}$ is called the protograph matrix of the SC code, and it has the convolutional structure composed of $L$ replicas $\mathbf{R}_{i}^{\textup{g}}$ of size $(m+L)\gamma\times\kappa$, where
@@ III-A Probabilistic Metrics
$\sum_{i=0}^{m}\sum_{j=0}^{M-1}p_{i,j}=1,$ (32)
and each $p_{i,j}$ specifies the probability that a $1$ in $\mathbf{H}_{\textup{MD}}^{\textup{g}}$ is assigned to the $i^{\textup{th}}$ component of the $j^{\textup{th}}$ auxiliary matrix. Thus, $\mathbf{P}$ is referred to as the (joint) probability-distribution matrix. Recall that $\mathbf{X}_{0}\triangleq\mathbf{H}_{\textup{SC}}^{\prime}$. The two-variable coupling polynomial of an MD-SC code associated with the probability-distribution matrix $\mathbf{P}$ is
@@
Denote by $N_{8}(\mathbf{p}^{\textup{con}})$ the expected number of cycle-$8$ candidates in the MD protograph. Then,
@@
where $w_{1}=\binom{\gamma}{2}\binom{\kappa}{2}$, $w_{2}=3\binom{\gamma}{2}\binom{\kappa}{3}+3\binom{\gamma}{3}\binom{\kappa}{2}$, $w_{3}=18\binom{\gamma}{3}\binom{\kappa}{3}$, $w_{4}=6\binom{\gamma}{2}\binom{\kappa}{4}+6\binom{\gamma}{4}\binom{\kappa}{2}+36\binom{\gamma}{3}\binom{\kappa}{4}+36\binom{\gamma}{4}\binom{\kappa}{3}+24\binom{\gamma}{4}\binom{\kappa}{4}$ if $\gamma\geq 4$, and $w_{4}=6\binom{\gamma}{2}\binom{\kappa}{4}+36\binom{\gamma}{3}\binom{\kappa}{4}$ if $\gamma=3$, where $\kappa\gg 4$ in practice.
@@
Observe that there are nine protograph patterns that can produce cycles-$8$ after lifting [15] (see Fig. 1). Following the logic in Theorem 1, [8, Theorem 1], and their proofs, we can derive the result above. ∎
@@
In this paper, we do not discuss cycles-$4$ in the analysis since removing all of them via informed lifting is easy. We also consider prime $z$, which implies that protograph cycles-$4$ cannot produce cycles-$8$ after lifting. Therefore, we take $w_{1}=0$ in the version of Algorithm 1 that handles cycles-$8$ in Section IV to obtain the probability-distribution matrix.
@@ III-B Gradient-Descent MD-SC Solution Form
The probabilities $p_{i}^{*}$’s of the vector $\mathbf{p}^{*}$ in the constraints of Lemma 1 and 2 introduced below are set to a locally-optimal edge distribution for the underlying SC code obtained as in [8]. Since they add up to $1$, (32) is automatically satisfied. We chose this approach in order to apply MD relocations to the best underlying SC code. These probabilities are used as an input to Algorithm 1 in Section IV.
@@
for all $i\in\{0,1,\dots,m\}$, it is necessary that the following equations hold for some $c_{i}\in\mathbb{R}$:
@@
for all $i\in\{0,1,\dots,m\}$, it is necessary that the following equations hold for some $c_{i}\in\mathbb{R}$:
$\sum_{M|b}\Big{\{}w_{1}\left[4f^{2}(X,Y)f(X^{-1},Y^{-1})\right]_{i,b+j}$
@@
$=\sum_{M|b}\nabla_{\mathbf{p}^{\textup{con}}}\Big{\{}w_{1}\left[f^{2}(X,Y)f^{2}(X^{-1},Y^{-1})\right]_{0,b}$
@@
$=\sum_{M|b}\Big{\{}w_{1}\left[4f^{2}(X,Y)f(X^{-1},Y^{-1})\mathbf{v}_{1}^{\textup{con}}\right]_{0,b}$
@@ III-C Outcomes Forecasted
where $L>\chi$ and $\chi$ is the maximum number of consecutive SC replicas a cycle of interest can span ($\kappa\gg 4$ in practice).
@@
where $s_{\textup{min}}$ and $s_{\textup{max}}$ are the minimum and maximum spans, in terms of replicas, of a cycle-$6$ in $\mathbf{H}_{\textup{SC}}$. Note that $s_{\textup{min}}=1$ and $s_{\textup{max}}=\chi=m+1$. By assumption, $L\gg m$, and thus we can take $E_{6}(\mathbf{p}^{\textup{con}})=E_{\textup{obj}}\cdot\frac{(2L-m)}{2}\cdot M$ via averaging $L$ and $L-m$ as an estimate, which gives (19).
@@
where $s_{\textup{min}}$ and $s_{\textup{max}}$ are the minimum and maximum spans of a cycle-$8$ in $\mathbf{H}_{\textup{SC}}$. Note that $s_{\textup{min}}=1$ and $s_{\textup{max}}=\chi=2m+1$ due to the existence of structures $P_{6}$ and $P_{7}$ (only $P_{7}$) in case $\gamma\geq 4$ ($\gamma=3$). By assumption, $L\gg m$, and thus we can take $E_{8}(\mathbf{p}^{\textup{con}})=N_{8}(\mathbf{p}^{\textup{con}})\cdot(L-m)\cdot M$ via averaging $L$ and $L-2m$ as an estimate, which gives (20). ∎
@@ IV-A Multi-Dimensional Gradient-Descent Distributor
1:Inputs: $\gamma,\kappa,L,m,M$: parameters of the MD code, $\mathbf{p}^{*}$: locally-optimal edge distribution of an SC code with $\gamma,\kappa,m$ obtained by [8, Algorithm 1], $\mathcal{T}$: MD density, and $\epsilon,\alpha$: accuracy and step size of gradient-descent.
@@
    Then, it corrects its rows by making the negative entries 00 and by updating the positive entries in order not to change the sum of each row via
@@ IV-B Probabilistic Design of GD-MD Codes
7:For each circulant $\mathcal{C}$, determine the number of active cycles which involve $\mathcal{C}$ from the pruned cycle list and select the circulant(s) with the maximum number. If multiple circulants, choose one randomly.
8:For each $i\in\{0,1,\dots,M-1\}$, determine the number of cycles involving $\mathcal{C}$, determined in Step 4, that become inactive after relocating $\mathcal{C}$ to $\mathbf{X}_{i}$, based on the relocation condition in (III-A) (see [23]).
@@ V Experimental Results and Comparisons
In this section, all GD-MD codes are constructed by FL-AO Algorithm with initial relocations based on the probability-distribution matrix obtained by MD-GRADE Algorithm. We use SC codes constructed via discrete optimization (the OO-CPO approach) [13] and via gradient-descent (the GRADE-AO approach) [8], as well as a topologically-coupled (TC) code from [8] as the underlying SC codes. Iterations here are always FL-AO iterations, and rates are always design rates. We first provide our results, then discuss the takeaways, and finally compare our codes with those in the literature.
@@
In the following, SC Code x is the underlying SC code of GD-MD Code x.y, and SC Code x.y is the 1D counterpart of GD-MD Code x.y, i.e, SC Code x.y has the same replica $\mathbf{R}_{1}$ as SC Code x and has the same length as GD-MD Code x.y. Our code parameters are as follows.
@@
Below are some takeaways from these results:
@@
    Since the partitioning and lifting matrices of the underlying SC codes are designed in a way that minimizes the number of cycles (of length $=$ girth), the final number of cycles we obtain after FL-AO Algorithm is lower than the upper bound on the expected number of cycles obtained in Theorem 3. For GD-MD Code 3.1, Theorem 3 gives a good indicator of the outcome of FL-AO Algorithm, where the (rounded) expected number of cycles-$8$ is $228{,}070$ and the actual final number of cycles-$8$ is $239{,}752$. Similarly, for GD-MD Code 2.2, the expected number of cycles-$8$ is $2{,}222{},235$, its upper bound is $3{,}584{,}250$, and the actual final number of cycles-$8$ is $2{,}768{,}485$, falling in between. However, the expected number of cycles-$8$ of GD-MD Code 4.1 is $396{,}434$ and the actual number of cycles-$8$ obtained is $112{,}959$, illustrating the strength of FL-AO Algorithm.
@@ VI Conclusion
In this work, we developed a probabilistic framework based on the GD algorithm to design MD-SC codes. We attempt to minimize the number of short cycles (of length $=$ girth) in the Tanner graph of such codes. We defined a probability-distribution matrix that characterizes the MD code design. We expressed the expected number of detrimental cycles in the GD-MD code in terms of probabilities in the distribution matrix. With this expected number in hand, we are able to answer two major questions, which will be of value to the coding theory community. The first is about an estimate of the percentage relocations required to remove all instances of a cycle in case it can be removed entirely. The second is about the best that can be done regarding a cycle given a pre-specified maximum relocation percentage, imposed by decoding latency requirements. Our MD-GRADE algorithm provides us with a locally-optimal distribution matrix that guides the FL-AO by reducing its search space. We demonstrated that, when fed with a locally-optimal probability-distribution matrix, the FL-AO can converge on excellent finite-length MD-SC designs (GD-MD codes) in a notably fast manner. Our GD-MD codes achieve notable reduction in the number of short cycles. Future work includes extending the analysis to objects that are more advanced than short cycles.

### v4

@@ I Introduction
The finite-length performance advantage of SC codes results from the additional degrees of freedom they offer the code designer via partitioning. Those additional degrees of freedom grow with the memory of the SC code. There are various results on high performance SC code designs with low memories in the literature [7, 11, 12]. Some of these designs are optimal with respect to the number of detrimental objects [13]; others demonstrated notable performance gains in data storage and transmission systems [14, 15]. Optimal solutions are notorious to achieve for SC codes with high memories because of the remarkable complexity growth. Therefore, a recent idea introduced probabilistic design of SC codes with high memories, which is based on the gradient-descent (GD or GRADE) algorithm, to find locally-optimal solutions with respect to the number of detrimental objects [8].
@@ II Preliminaries
$\mathbf{H}_{\textup{SC}}$ is obtained from some binary matrix $\mathbf{H}^{\textup{g}}_{\textup{SC}}$ by replacing each nonzero (zero) entry in it by a circulant (zero) $z\times z$ matrix, $z\in\mathbb{N}$. Throughout the paper, the notation “g” in the superscript of any matrix refers to its protograph matrix, where each circulant matrix is replaced by $1$ and each zero matrix is replaced by 00. $\mathbf{H}^{\textup{g}}_{\textup{SC}}$ is called the protograph matrix of the SC code, and it has the convolutional structure composed of $L$ replicas $\mathbf{R}_{i}^{\textup{g}}$ of size $(m+L)\gamma\times\kappa$, where
@@ III-A Probabilistic Metrics
$P_{6}(\mathbf{p}^{\textup{con}})=\mathbb{P}\left[\sum_{k=1}^{3}[\mathbf{K}(i_{k},j_{k})-\mathbf{K}(i_{k},j_{k+1})]=0,\sum_{k=1}^{3}[\mathbf{M}({i_{k},j_{k}})-\mathbf{M}({i_{k},j_{k+1}})]\equiv 0\textup{ (mod $M$)}\right]$
$=$ $\sum\limits_{\begin{subarray}{c}\sum\nolimits_{k=1}^{3}(x_{k}-y_{k})=0,\\ \sum\nolimits_{k=1}^{3}(u_{k}-v_{k})\equiv 0\textup{ (mod $M$)}\end{subarray}}\prod_{k=1}^{3}\mathbb{P}\left[\mathbf{K}(i_{k},j_{k})=x_{k},\mathbf{M}({i_{k},j_{k}})=u_{k},\mathbf{K}(i_{k},j_{k+1})=y_{k},\mathbf{M}({i_{k},j_{k+1}})=v_{k}\right]$
$=$ $\sum\limits_{\begin{subarray}{c}\sum\nolimits_{k=1}^{3}(x_{k}-y_{k})=0,\\ \sum\nolimits_{k=1}^{3}(u_{k}-v_{k})\equiv 0\textup{ (mod $M$)}\end{subarray}}p_{x_{1},u_{1}}p_{x_{2},u_{2}}p_{x_{3},u_{3}}p_{y_{1},v_{1}}p_{y_{2},v_{2}}p_{y_{3},v_{3}}$
$=$ $\sum_{M|b}\left[\sum\limits_{\begin{subarray}{c}x_{k},y_{k}\in\{0,1,\dots,m\},\\ u_{k},v_{k}\in\{0,1,\dots,M-1\}\end{subarray}}p_{x_{1},u_{1}}p_{x_{2},u_{2}}p_{x_{3},u_{3}}p_{y_{1},v_{1}}p_{y_{2},v_{2}}p_{y_{3},v_{3}}X^{x_{1}+x_{2}+x_{3}-y_{1}-y_{2}-y_{3}}Y^{u_{1}+u_{2}+u_{3}-v_{1}-v_{2}-v_{3}}\right]_{0,b}$
$=$ $\sum_{M|b}\left[f^{3}(X,Y)f^{3}(X^{-1},Y^{-1})\right]_{0,b}.$ (9)
@@
$\sum_{i=0}^{m}\sum_{j=0}^{M-1}p_{i,j}=1$ (5)
and each $p_{i,j}\in[0,1]$ specifies the probability that a $1$ in $\mathbf{H}_{\textup{MD}}^{\textup{g}}$ is assigned to the $i^{\textup{th}}$ component of the $j^{\textup{th}}$ auxiliary matrix. Thus, $\mathbf{P}$ is referred to as the (joint) probability-distribution matrix. Recall that $\mathbf{X}_{0}\triangleq\mathbf{H}_{\textup{SC}}^{\prime}$. The two-variable coupling polynomial of an MD-SC code associated with the probability-distribution matrix $\mathbf{P}$ is
@@
Denote by $N_{8}(\mathbf{p}^{\textup{con}})$ the expected number of cycle-$8$ candidates (see Fig. 1) in the MD protograph. Then,
@@
where $w_{1}=\binom{\gamma}{2}\binom{\kappa}{2}$, $w_{2}=3\binom{\gamma}{2}\binom{\kappa}{3}+3\binom{\gamma}{3}\binom{\kappa}{2}$, $w_{3}=18\binom{\gamma}{3}\binom{\kappa}{3}$, $w_{4}=6\binom{\gamma}{2}\binom{\kappa}{4}+6\binom{\gamma}{4}\binom{\kappa}{2}+36\binom{\gamma}{3}\binom{\kappa}{4}+36\binom{\gamma}{4}\binom{\kappa}{3}+72\binom{\gamma}{4}\binom{\kappa}{4}$ if $\gamma\geq 4$, and $w_{4}=6\binom{\gamma}{2}\binom{\kappa}{4}+36\binom{\gamma}{3}\binom{\kappa}{4}$ if $\gamma=3$, where $\kappa\geq 4$.
@@
Observe that there are nine protograph patterns that can produce cycles-$8$ after lifting [15] (see Fig. 1). Following the logic in Theorem 1, [8, Theorem 1], and their proofs, we can derive the result above (see also [26, Chapter 5]). ∎
@@
In this paper, we do not discuss cycles-$4$ in the analysis since removing all of them via informed lifting is easy. We also consider prime $z$, which implies that protograph cycles-$4$ cannot produce cycles-$8$ after lifting. Thus, and since relocations come after lifting in the FL code construction, we take $w_{1}=0$ in the version of Algorithm 1 that handles cycles-$8$ in Section IV to obtain the probability-distribution matrix.
@@ III-B Gradient-Descent MD-SC Solution Form
The vector $\mathbf{p}^{*}$ of the probabilities $p_{i}^{*}$’s in the constraints of Lemma 1 and Lemma 2 introduced below are set to a locally-optimal edge distribution for the underlying SC code obtained as in [8]. Since they add up to $1$, (5) is automatically satisfied. We chose this approach in order to apply MD relocations to the best underlying SC code. These probabilities are used as an input to Algorithm 1 in Section IV.
@@
for all $i\in\{0,1,\dots,m\}$, where each $p_{i,j}\in[0,1]$, it is necessary that the following equations hold for some $c_{i}\in\mathbb{R}$:
@@
for all $i\in\{0,1,\dots,m\}$, where each $p_{i,j}\in[0,1]$, it is necessary that the following equations hold for some $c_{i}\in\mathbb{R}$:
$\sum_{M|b}\Big{\{}w_{1}\left[4f^{2}(X^{2},Y^{2})f(X^{-2},Y^{-2})\right]_{2i,b+2j}$
@@
$=\sum_{M|b}\nabla_{\mathbf{p}^{\textup{con}}}\Big{\{}w_{1}\left[f^{2}(X^{2},Y^{2})f^{2}(X^{-2},Y^{-2})\right]_{0,b}$
@@
$=\sum_{M|b}\Big{\{}w_{1}\left[4f^{2}(X^{2},Y^{2})f(X^{-2},Y^{-2})\mathbf{v}_{2}^{\textup{con}}\right]_{0,b}$
@@ III-C Outcomes Forecasted
where $L>\chi$ and $\chi$ is the maximum number of consecutive SC replicas a cycle of interest can span ($\kappa\geq 4$).
@@
where $s_{\textup{min}}$ and $s_{\textup{max}}$ are the minimum and maximum spans, in terms of replicas, of a cycle-$6$ in $\mathbf{H}_{\textup{SC}}$. Note that $s_{\textup{min}}=1$ and $s_{\textup{max}}=\chi=m+1$. By assumption, $L\gg m$, and thus we can take $E_{6}(\mathbf{p}^{\textup{con}})\approx E_{\textup{obj}}\cdot\frac{(2L-m)}{2}\cdot M$ via averaging $L$ and $L-m$ as an estimate, which gives (19).
@@
where $s_{\textup{min}}$ and $s_{\textup{max}}$ are the minimum and maximum spans of a cycle-$8$ in $\mathbf{H}_{\textup{SC}}$. Note that $s_{\textup{min}}=1$ and $s_{\textup{max}}=\chi=2m+1$ due to the existence of structures $P_{6}$ and $P_{7}$ (only $P_{7}$) in case $\gamma\geq 4$ ($\gamma=3$). By assumption, $L\gg m$, and thus we can take $E_{8}(\mathbf{p}^{\textup{con}})\approx N_{8}(\mathbf{p}^{\textup{con}})\cdot(L-m)\cdot M$ via averaging $L$ and $L-2m$ as an estimate, which gives (20). ∎
@@ IV-A Multi-Dimensional Gradient-Descent Distributor
1:Inputs: $\gamma,\kappa,L,m,M$: parameters of the MD code, $\mathbf{p}^{*}$: locally-optimal edge distribution of an SC code with $\gamma,\kappa,m$ obtained by [8, Algorithm 1], $\mathcal{T}$: MD density, and $\epsilon,\alpha$: accuracy and step size of gradient descent.
@@
    Summations are for non-updated values. Then, it corrects its rows by making the negative entries 00 and by updating the positive entries in order not to change the sum of each row via
@@ IV-B Probabilistic Design of GD-MD Codes
7:For each circulant, determine the number of active cycles which involve it from the pruned cycle list and select the circulant(s) with the maximum number. If multiple circulants, choose one randomly. Let $\mathcal{C}$ denote the selected circulant.
8:For each $i\in\{0,1,\dots,M-1\}$, determine the number of cycles involving $\mathcal{C}$ that become inactive after relocating $\mathcal{C}$ to $\mathbf{X}_{i}$, based on the relocation condition in (III-A) (see [23]).
@@ V Experimental Results and Comparisons
In this section, all GD-MD codes are constructed by FL-AO Algorithm with initial relocations based on the probability-distribution matrix obtained by MD-GRADE Algorithm. We use SC codes constructed via discrete optimization (the OO-CPO approach) [13] and via gradient descent (the GRADE-AO approach) [8], as well as a topologically-coupled (TC) code from [8] as the underlying SC codes. Iterations here are always FL-AO iterations, and rates are always design rates. We first provide our results, then discuss the takeaways, and finally compare our codes with those in the literature.
@@
In the following, SC Code x is the underlying SC code of GD-MD Code x.y, and SC Code x.y is the 1D counterpart of GD-MD Code x.y, i.e, SC Code x.y has the same replica $\mathbf{R}_{1}$ as SC Code x and has the same length as GD-MD Code x.y. Our code parameters and results are as follows.
@@
Below are some takeaways from these results, which are also summarized in Table II:
@@
    Since the partitioning and lifting matrices of the underlying SC codes are designed in a way that minimizes the number of cycles (of length $=$ girth), the final number of cycles we obtain after FL-AO Algorithm is lower than the upper bound on the expected number of cycles obtained in Theorem 3. For GD-MD Code 3.1, Theorem 3 gives a good indicator of the outcome of FL-AO Algorithm, where the (rounded) expected number of cycles-$8$ is $228{,}070$ and the actual final number of cycles-$8$ is $239{,}752$. Similarly, for GD-MD Code 2.2, the expected number of cycles-$8$ is $2{,}681{,}009$, its upper bound is $4{,}324{,}208$, and the actual final number of cycles-$8$ is $2{,}768{,}485$, falling in between, but quite close to the estimate provided by the aforementioned expected number. This observation further demonstrates the value of Theorem 3. However, the expected number of cycles-$8$ of GD-MD Code 4.1 is $396{,}434$ and the actual number of cycles-$8$ obtained is $112{,}959$, illustrating the strength of FL-AO Algorithm.
@@ VI Conclusion
In this work, we developed a probabilistic framework based on the GD algorithm to design MD-SC codes. We attempt to minimize the number of short cycles (of length $=$ girth) in the Tanner graph of such codes. We defined a probability-distribution matrix that characterizes the MD code design. We expressed the expected number of detrimental cycles in the GD-MD code in terms of probabilities in the distribution matrix. With this expected number in hand, we are able to answer two major questions, which will be of value to the coding theory community. The first is about an estimate of the percentage relocations required to remove all instances of a cycle in case it can be removed entirely. The second is about the best that can be done regarding a cycle given a pre-specified maximum relocation percentage, imposed by decoding latency requirements. Our MD-GRADE algorithm provides us with a locally-optimal distribution matrix that guides the FL-AO by reducing its search space. We demonstrated that, when fed with a locally-optimal probability-distribution matrix, the FL-AO can converge on excellent finite-length MD-SC designs (GD-MD codes) in a notably fast manner. Our GD-MD codes achieve notable reduction in the number of short cycles. Future work includes simulating the codes and extending the analysis to objects that are more advanced than short cycles.
VII Appendix